shift summary eve 10/16/99

Patrick Toale (toale@cuhep.Colorado.EDU)
Sat, 16 Oct 1999 23:23:21 -0600 (MDT)

Shift Summary for eve shift, 16/Oct/99

On Shift: Pat T., Taku

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 8.5
Hours of E799 data (special runs not included): 7.0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: changing run/tapes, pipeline errors

========================================================================
Runs ended this shift:
Run 14850 - 22.4M events - BEAM_799

Tapes brought to FCC this shift: UP*85-87,UPS029
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
E799 datataking, L3 investigation - hoped to open champagne but were
held back by others

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14850 15:30 Inherit shift.

15:52 Beam lost. Est 15 min.

16:09 Dead 'scaler_check' process. Paged Ashkan, got
it restarted.

16:55 Tapes full. Stop run, run does not stop properly.
Fresh start, pipeline errors. Hard reboot crate 10.

14852 17:37 Start run.

20:30 Jason L. deals with ~95% full onsplit disks.

220 22:30 Laser illum/ped problems with ch 906. More
pipeline problems, reboot crates 11,12,15,16.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

dc#summary : planes 10,11 eff. is sad.

Thanks,
Pat Toale
--------------------------------------------------------------
Graduate Research Assistant
Department of Physics University of Colorado at Boulder
Office: F-231 Phone: (303) 492-4363
E-Mail: toale@cuhep.colorado.edu